If you have a piece of pipe that's 180 inches. It is cut into 2 pieces. The largest piece is 5 times longer than the smaller piece what are the lengths of the 2 pieces of pipe?
Found 2 solutions by Fombitz, Alan3354:
Answer by Fombitz(32388)
Answer by Alan3354(69443)

If you have a piece of pipe that's 180 inches. It is cut into 2 pieces. The largest piece is 5 times longer than the smaller piece what are the lengths of the 2 pieces of pipe?
============================
5 times longer = 6 times as long.
S = length of small piece.
---
S + 6S = 180
S = 180/7 inches
Longer piece = 1080/7 inches.
===================
PS If "5 times longer" = 5 times as long, then what is 1 time longer, or 100% longer?
Is it the same length?